please help me solve this equation find the mean,median, mode,sd of 1,3,2,5,4,2,4
Calculate z150 mean is 100 sd= 15

Answer by stanbon(75887)

find the mean,median, mode,sd of 1,3,2,5,4,2,4
mean = (sum of the numbers)/7 = 3
median is the middle number = 4
mode is the most frequent number = 4
------------------------------------------
Calculate z150 mean is 100 sd= 15
z(150) = (150-100)/15 = 50/15 = 10/3 = 3.33
-----------------------
Cheers,
Stan H.
---------------